Tuesday 06/28/11

9:30 A-7:00P, 7:40-8:40, 10:00P- 11:10P

  • Data alignment
  • Flip sna[D] and hb[D] flies into new bottles, keep at 25C.
  • Compare hb data by image type.  confirm fairly dynamic expression prevents easy super-position of patterns during cc13.
  • February’s data format doesn’t fit into todays plotting commands.  Running new analysis of het data from February. Going to be slow to get all this through the pipeline.   3 parallel counting scripts pretty much maxes out computer resources.
  • Consolidate AP-toll, AP-twi1, AP-twi2, GD7 and toll[10b] embryos.
  • Controls: comparing separate acquisitions of same slide in different y-hb-BAC contexts to check for reproducibility.
  • Data analysis issues: low exposure images with only 2 consec layers in dot-finding produce too much noise and background in measurement.   Rerun s03_MP02 data.
  • Rerunning hets MP09_c and MP09_d at .03, .01 with revised Unsupervised_DotFinding.m
  • Running G4B_LacZ analysis
This entry was posted in Summaries and tagged , , , , . Bookmark the permalink.

Leave a Reply

Your email address will not be published. Required fields are marked *